OpenStudy (anonymous):

Determine whether the sequence: ln(2n^2 +1) - ln(n^2 +1) converges or diverges. If the sequence converges, find the limit. I tried expressing the sequence as: ln(2n^2 +1/n^2 +1), but still no luck. The answer in the back of the book is ln(2) I am not sure how to manipulate the equation in order to take the limit. Can I solve this using squeeze theorum?

OpenStudy (anonymous):

i believe this should do the trick: \[\ln(2n^{2}+1)-\ln(n^{2}+1) = \ln\frac{2n^{2}+1}{n^{2}+1}\] The limit of the fraction will be 2, so basically as n goes to infinity, the sequence converges to ln 2

OpenStudy (anonymous):

if you need me to clear up anything, just ask :)

OpenStudy (anonymous):

wouldn't the fraction end up being 2(∞)^2 +1/(∞)^2+1 be an indeterminate form in which I would need to apply l'hospitals rule?

OpenStudy (anonymous):

ohhhh I see what you meant!

OpenStudy (anonymous):

thanks now I get it

OpenStudy (anonymous):

sweet :)